Solve: |2x + 4| = 12
I thought the answer was 12. Is that right?
Thank you very much

Answer by Ave(106)

Write two equations without the absolute value bars
2x+4=12 and -(2x+4)=12 solve both
2x=8 -2x=16
x=4 x=-8
{-8, 4}